# Superexchanges and Charge Transfer in the La3Ni2O7 Thin Films

### v1

@@ Abstract
To shed light on the effective low-energy model study that neglect oxygen degree of freedom, we group the Ni-$3d$ and O-$p$ orbitals into two classes: the out-of-plane (OP) correlated orbitals, including $d_{3z^{2}-r^{2}}$, $p_{z}$ and $p_{z}^{\prime}$ and the in-plane (IP) correlated orbitals, comprising $d_{x^{2}-y^{2}}$ and $p_{x}/p_{y}$ orbitals. And analyze their carrier concentration under hole/electron doping. The results, shown in Figure 4 (d), indicated that the additional holes doped into the LNO film are distributed nearly equally between the IP and OP orbitals, with a slight excess of holes in the IP orbitals. In the case of electron doping, the ratio of doped electrons in the IP to the OP orbitals is approximately 3:1 until the system reaches half-filling. This contrasts with the 1:1 ratio observed in hole doping, which suggests an asymmetric doping-dependent superconducting phase diagram. In experiments, Sr-doping introduce extra holes into the system [66], while possible Ce-doping can introduce extra electrons into the pristine system.
@@
In comparison to bulk LNO, we find that under the same chemical potential variation, the doping variable of LNO film is approximately twice that of bulk LNO [15]. This suggests that LNO films exhibit enhanced charge transfer capabilities compared to the bulk LNO under in-plane biaxial compression, meaning the LNO thin film possesses a smaller charge transfer gap $\Delta$ [72]. This finding aligns with recent spectroscopic experimental results [73], which propose that two types of Zhang–Rice singlets exist in the LNO film and that the fundamental classification of in-plane charge transfer properties is identical to that of the cuprate superconductors.
@@
Furthermore, it is discovered that the presence of Hund’s coupling $J_{H}$ in LNO films under half-filling strongly suppresses orbital polarization, thereby stabilizing a high-spin $S=1$ state. This phenomenon is illustrated by the gray and red dashed lines in Figure 4 (b), which correspond to the Hund’s first law. In contrast, in the pristine state, the $J_{H}$ coupling is significantly localized, and holes still preferentially occupy the in-plane orbitals, as indicated by the gray and red solid lines in Figure 4 (b).
@@
Discussion and conclusion.- The observation of superconductivity in La3Ni2O7 thin films at ambient pressure [49, 50] has spurred proposals of several effective interacting models to investigate the pairing symmetry [74, 75, 51, 52], building on insights from ab initio calculations [47, 48, 54, 74, 76]. These effective models primarily focus on parameterizing the magnetic correlations between Ni $d$-orbitals through direct exchange couplings, while the contribution of oxygen degrees of freedom is neglected. In contrast our 11-band Hubbard model, which incorporates both Ni-$3d$ and O-$2p$ orbitals, provides a more comprehensive picture by quantifying the relative strengths of superexchange couplings between distinct Ni-$3d$ orbitals. Crucially, we identify two spin-singlet bands exhibiting markedly different hole concentrations ($n_{h}$) and correlation strengths. We observe the significant reduction of IT magnetic exchange coupling, which makes the IR and IT coupling strengths comparable, in sharp contrast to the IT-dominated bulk material. Furthermore, we observe a pronounced asymmetry in charge distribution upon doping. It should be noted that some studies suggest Hund’s coupling $J_{H}$ can transfer antiferromagnetic correlation from $d_{3z^{2}-r^{2}}$ orbital to $d_{x^{2}-y^{2}}$ orbital [21, 20, 19, 77]. Based on our results, we find that a large $J_{H}$ does indeed enhance the inter-layer $d_{x^{2}-y^{2}}$ correlation. However, the large $J_{H}$ does not significantly impact the carriers distribution in the $e_{g}$ orbitals. Our analysis of both spin-spin correlations and charge-transfer properties consistently indicates that the inter-layer $Ni-3d_{3z^{2}-r^{2}}$ orbital antiferromagnetic exchange interaction remains an important factor that primarily determines the nature of the bilayer nickelate superconducting pairing. And the contribution of more itinerant Ni-$d_{x^{2}-y^{2}}$ orbital in to both strange-metal (SM) and superconducting (SC) states remains complex and requires further investigations.
@@
Acknowledgements.- We thank the useful discussion with Zhihui Luo and Wenyuan Qiu. This project was supported by NSFC-12494591, NSFC-92165204, NSFC-92565303, NKRDPC-2022YFA1402802, Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices (2022B1212010008), Research Center for Magnetoelectric Physics of Guangdong Province (2024B0303390001), and Guangdong Provincial Quantum Science Strategic Initiative (GDZX2401010).

### v2

@@ Abstract
To shed light on low-energy effective models integrating out oxygen degrees of freedom, we group the Ni-$3d$ and O-$2p$ orbitals into two categories: (i) the out-of-plane (OP) correlated orbitals, including $d_{3z^{2}-r^{2}}$, $p_{z}$ and $p_{z}^{\prime}$ and (ii) in-plane (IP) correlated orbitals comprising $d_{x^{2}-y^{2}}$ and $p_{x}/p_{y}$ orbitals. As shown in Figure 4 (d), additional holes doped into LNO film distribute nearly equally between IP and OP orbitals. In contrast, for electron doping around $\mu_{e}\sim 0$, the doped carriers favor IP orbitals over OP orbitals by a ratio of approximately 3:1. This stark asymmetry between electron and hole doping (1:1 vs. 3:1) suggests a dopant-type dependent superconducting phase diagram. This is particularly relevant to experiment, where Sr-doping introduces holes [68], while Thorium (Th) doping could introduce electrons.
@@
The data in Figure 4(d) also show a significant difference in carrier response: under identical chemical potential changes, the hole variation in LNO films is about twofold greater than in the bulk [15]. This indicates that LNO thin films possess enhanced charge transfer capabilities and, consequently, a smaller charge transfer gap $\Delta$ [74]. Our conclusion aligns with recent spectroscopic experiments [75], which propose the existence of two types of Zhang–Rice singlets in LNO films and suggest the in-plane charge transfer properties are analogous to those of cuprate superconductors.
@@
Finally, we find that the presence of Hund’s coupling $J_{H}$ at half-filling strongly in general suppresses orbital polarization, hence tends to stabilize a high-spin $S=1$ state, as illustrated by the dashed lines in Figure 4 (b), which follows the Hund’s first law. However, for the pristine configuration ($\mu=0$), the influence of the $J_{H}$ coupling is significantly weaken. In this case, holes retain a preferential occupation of the in-plane orbitals, as indicated by the solid lines in Figure 4 (b).
@@
Discussion and conclusion.- The observation of superconductivity in La3Ni2O7 thin films at ambient pressure [51, 52] has spurred proposals of several effective interacting models for investigating various physical properties [76, 77, 53, 54], building on insights from ab initio calculations [49, 50, 56, 76, 78]. These effective models primarily focus on parameterizing the magnetic correlations based on direct exchange couplings between nickel’s $d$-orbitals, while the contribution of oxygen degrees of freedom is neglected. In this work, the 11-band dpd-p Hubbard model, which incorporates both Ni-$3d$ and O-$2p$ orbitals, provides a more comprehensive picture for understanding the relative strengths of superexchange couplings and distributions of charge carriers in La3Ni2O7 thin films. Crucially, we resolve properties of in-plane and out-of-plane bands exhibiting remarkable different hole concentrations ($n_{h}$) and correlation strengths. We find that IR and IT magnetic coupling strengths are more or less comparable, in sharp contrast to the IT coupling - dominated bulk materials. Furthermore, we observe a pronounced asymmetry in charge distribution upon hole/electron doping. In summary, our unified analysis on both magnetic correlations and charge-transfer properties consistently identifies the interlayer $Ni-3d_{3z^{2}-r^{2}}$ orbital couplings as the dominant factor determining the nature of pairing in the bilayer nickelate films. The contribution of the more itinerant Ni-$d_{x^{2}-y^{2}}$ orbitals in superconducting or normal state, however, appear more subtle and complex, warranting further dedicated investigations.
@@
Acknowledgements.- We thank the useful discussion with Zhihui Luo and Wenyuan Qiu. This project was supported by NSFC-12494591, NSFC-12494594, NSFC-92165204, NSFC-92565303, NKRDPC-2022YFA1402802, Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices (2022B1212010008), Research Center for Magnetoelectric Physics of Guangdong Province (2024B0303390001), and Guangdong Provincial Quantum Science Strategic Initiative (GDZX2401010).